# Matchings in hypercubes extend to long cycles

### v1

@@ Abstract.
This work was supported by Czech Science Foundation grant GA 22-15272S
@@ 1. Introduction
Figure 1. Cycles and matchings in the hypercube $Q_{4}$: (a) a Hamilton cycle, where the vertex labels omit curly brackets and commas for conciseness; (b) a perfect matching $M$ and an extension of $M$ to a Hamilton cycle; (c) a maximal, but not perfect matching $M^{\prime}$; (d) an extension of $M^{\prime}$ to a cycle factor with two cycles; (e) an extension of $M^{\prime}$ to a (non-Hamilton) cycle; (f) an extension of $M^{\prime}$ to a Hamilton cycle (obtained by joining the two cycles in (d)).
@@
In fact, the strengthening to consider $K(Q_{d})$ instead of $Q_{d}$ is the key idea to Fink’s proof of Theorem 2, as it makes the induction hypothesis stronger and thus more flexible. This construction actually shows that every matching of $K(Q_{d})$ can be extended to at least $2^{2^{d-4}}$ distinct Hamilton cycles. This proof technique has been exploited in several subsequent papers (see e.g. [Gre09, GNŠ18, ${}^{+}$15]), and we will also use it heavily in our arguments.
@@ 1.1. Our results
Furthermore, we can give some guarantee about the length of the cycle obtained from our construction, namely that it visits a constant fraction of the vertex set $V(Q_{d})$.
@@
To prove Theorem 5 inductively, we will need the following theorem, which yields a cycle that extends a given matching but also avoids one ‘forbidden’ vertex $z$, i.e., avoiding $z$ is an additional constraint imposed on the cycle. Using that $Q_{d}$ is vertex-transitive, we may assume w.l.o.g. that the forbidden vertex is $z=\emptyset$. The following theorem requires some additional conditions, and to state them we need to introduce some notation; see Figure 3 (a): We split $Q_{d}$ in a direction $i\in[d]$ into subgraphs $Q^{i}_{0}$ and $Q^{i}_{1}$, where $Q^{i}_{1}$ is the subgraph induced by vertices of $Q_{d}$ containing $i$ and $Q^{i}_{0}$ is the subgraph induced by vertices of $Q_{d}$ not containing $i$. Note that $z\in V(Q^{i}_{0})$ for every $i\in[d]$. The set of edges of $Q_{d}$ that have one end vertex in $Q^{i}_{0}$ and the other in $Q^{i}_{1}$ is called a layer in direction $i$. This set splits into two equal-sized sets of edges, depending on the parity of their end vertices in $Q^{i}_{0}$, and we refer to each of these two sets of edges as a half-layer.
@@
Let $d\geq 5$. A matching $M$ of $K(Q_{d})$ that avoids $z=\emptyset$ can be extended to a cycle that avoids $z$ if and only it satisfies the following condition:
@@
We remark that the statement of Theorem 8 is false for $d=4$, and all counterexamples are listed in [GNŠ18].
@@ 1.2. Applications to Hamilton-laceability
Let $d\geq 5$, and let $x,y$ be two vertices of opposite parity in $Q_{d}$. A perfect matching $M$ of $B(Q_{d})$ can be extended to a Hamilton path with end vertices $x$ and $y$ if and only if $M\setminus\{xx^{M},yy^{M}\}\cup\{x^{M}y^{M}\}$ contains no half-layers.
@@
Let $d\geq 5$, and let $x,y$ be two vertices of opposite parity in $Q_{d}$. A perfect matching $M$ of $K(Q_{d})$ can be extended to a Hamilton path with end vertices $x$ and $y$ if and only if $M\setminus\{xx^{M},yy^{M}\}\cup\{x^{M}y^{M}\}$ contains no half-layers.
@@ 1.3. Avoiding and including other structures
Instead of forbidding edges, one may also consider the problem of forbidding certain vertices. To this end, Fink and Gregor [FG12] proved that for any set $F$ of at most $\binom{d}{2}-2$ vertices of $Q_{d}$, $d\geq 3$, there is a cycle of length at least $2^{d}-2|F|$ that avoids all vertices in $F$; see Figure 2 (d). Their result answer a conjecture by Castañeda and Gotchev [CG10], and the bound $\binom{d}{2}-2$ is best possible. Note that if we forbid some number of vertices of the same parity, then any extending cycle will also skip the same number of vertices of the opposite parity. This explains that we need to allow skipping $2|F|$ many vertices instead of only $|F|$.
@@ 1.4. Other restricted Gray codes
Hamilton cycles and paths in the hypercube are often referred to as Gray codes [Müt23, Sav97]. In this setting, we view the vertices of the hypercube as bitstrings of length $n$, given by the indicator vectors of the corresponding subsets. The well-known binary reflected Gray code, named after the Bell Labs researcher Frank Gray, is a particularly elegant construction of such a cycle that lends itself well to fast algorithmic computation, optimally so that each bit flipped is computed in constant time. There has been a large amount of work on Gray codes satisfying various other kinds of constraints, motivated by different applications:
@@
  • single-track Gray codes [Etz07, GMM22], where each bit follows the same cyclically shifted flipping pattern along the code.
@@ 1.5. Outline of this paper
In Section 2 we present some notations and terminology used throughout this paper, as well as some auxiliary lemmas needed later. In Sections 3 and 4 we present the inductive arguments for the proofs of Theorems 5 and 8, respectively. Section 5 contains the proofs of Theorems 6, 7, and 12.
@@ 2.1. Notation and definitions
For two sets $A$ and $B$, the symmetric difference of $A$ and $B$ is denoted by $A\mathbin{\triangle}B$. Recall that $Q_{d}$ has an edge between any two vertices $u$ and $v$ with $|u\mathbin{\triangle}v|=1$. For a vertex $u$ of $Q_{d}$ and $i\in[d]$, we define $u^{i}:=u\mathbin{\triangle}\{i\}$, and we refer to $i$ as the direction of the edge $uu^{i}$. We say that $u$ is even or odd, if $|u|$ is even or odd, respectively.
@@
For a graph $G$ and a subset of edges $M$ of $G$, we say that a vertex $u$ of $G$ is covered by $M$, if $u$ is end vertex of one of the edges of $M$. We write $V(M)$ for the set of all vertices covered by $M$. If $u$ is not covered by $M$ we say that it is avoided by $M$. Recall that for a vertex $u$ covered by an edge of a matching $M$, we write $u^{M}$ for the other end vertex of that edge of $M$. Furthermore, for $i\in[d]$ we use the abbreviation $u^{iM}:=(u^{i})^{M}$.
@@ 2.2. A cycle that avoids two vertices
If $C$ avoids $y$, then every edge of $C\setminus M$ has end vertices of opposite parity. Since $Q_{d}$ has the same number of even and odd vertices, it follows that the avoided vertices $x$ and $y$ also have opposite parity. On the other hand, if $C$ contains $y$, then $C\setminus M$ has edges whose end vertices have opposite parity, plus a path of length 2 whose two end vertices have opposite parity to the middle vertex $y$. It follows that $x$ must have the same parity as $y$. ∎
@@ 2.3. Properties of half-layers and quad-layers
To prove (i), let $i$ and $j$ be the directions of $L$ and $L^{\prime}$, respectively, and consider the decomposition of $Q_{d}$ into disjoint 4-cycles of the form $C=(u,u^{i},(u^{i})^{j},u^{j})$ for some $u\in V(Q_{d})$. From the two pairs of opposite edges in directions $i$ and $j$ in $C$, exactly one belongs to $L$ and one to $L^{\prime}$, respectively.
@@
The next lemma asserts that half-layers and quad-layers cannot exist simultaneously in several different directions.
@@
To prove (iii) we argue as follows. By Lemma 14 (i), two half-layers in different directions share $2^{d-2}$ vertices. Having two 2-near half-layers in different directions, each one uses $22^{d-2}-$ 2 of shared vertices to be matched in its direction (and not the other direction). However, $2(2^{d-2}-2)>2^{d-2}$ for $d\geq 5$, so two 2-near half-layers in different directions cannot exist for $d\geq 5$. ∎
@@
Let $d=5$ and let $M$ be a matching of $K(Q_{d})$ that contains a $z$-dangerous (near) quad-layer. Then there is a direction $i\in[d]$ such that $|M^{i}_{-}|\geq 3$ and $M^{i}_{0}$ contains no (near) half-layers of $Q^{i}_{0}$.
@@
Let $i_{1}$ be the direction of a $z$-dangerous (near) quad-layer, i.e., $|M^{i_{1}}_{-}|\geq 2^{d-3}-1=3$. If $M^{i_{1}}_{0}$ contains no $z$-dangerous (near) half-layers, then we are done. Otherwise, let $i_{2}$ be the direction of a $z$-dangerous (near) half-layer of $Q^{i_{1}}_{0}$, i.e., $|M^{i_{2}}|\geq 3$. If $M^{i_{2}}_{0}$ contains no $z$-dangerous (near) half-layers, then we are done. Otherwise, let $i_{3}$ be the direction of a $z$-dangerous (near) half-layer of $Q^{i_{2}}_{0}$, i.e., $|M^{i_{3}}|\geq 3$ etc. If this process terminates, then we are done. Otherwise, it loops in a cycle of length $\ell\geq 2$. As each of these $\ell$ many $z$-dangerous (near) quad-layers of $Q_{d}$ in direction $i_{j}$, $j\in\{1,\ldots,\ell\}$, covers at least $d-3=2$ neighbors of $z$ with an edge in direction $i_{j}$, we conclude that $\ell=2$. For $\ell=2$, we have a $z$-dangerous (near) half-layer of $Q^{i_{1}}_{0}$ in direction $i_{2}$ and a $z$-dangerous (near) half-layer of $Q^{i_{1}}_{0}$ in direction $i_{2}$, and each of them covers at least $d-3=2$ neighbors of $z$ in the subcube $(Q^{i_{1}}_{0})^{i_{2}}_{0}$ with an edge in direction $i_{2}$ and $i_{1}$, respectively. However, the vertex $z$ has only three neighbors in the 3-dimensional subcube $(Q^{i_{1}}_{0})^{i_{2}}_{0}$. ∎
@@
We first prove the statement for half-layers. We repeatedly match pairs of vertices in $A$, adding the matched pairs to $P$, thus decreasing the size of $A$ in each round, as follows: If $|A|\geq 6$, we match any two vertices of $A$ having the same parity, which does not create half-layers, and repeat. If $|A|=4$ we distinguish three cases. If all vertices in $A$ have the same parity, we match them arbitrarily, which does not create half-layers. If $A$ has two vertices of each parity, we match them in those pairs, and we are done as well. Otherwise we have one even and three odd vertices in $A$ or vice versa, and by symmetry it suffices to consider the first case. Let $a$ be the unique even vertex in $A$. If one of the three odd vertices in $A$ is not a neighbor of $a$ in $Q_{d}$, then we match $a$ to this vertex and the remaining two vertices to each other, which does not create half-layers. On the other hand, if all three odd vertices in $A$ are neighbors of $a$, then by Lemma 15 (i), for at most one of them adding an edge to $a$ may create a half-layer, leaving the other two odd vertices in $A$ to be matched with $a$ (and the remaining two vertices to each other), such that no half-layers are created.
Essentially the same proof works for near half-layers instead of half-layers, except the step where Lemma 15 (i) is invoked. This would require a statement as in Lemma 15 (iii), which however does not hold for $d=4$. Nonetheless, it is true that $M$ cannot contain 2-near half-layers in all three directions from $a$ to the three odd neighbors of $a$ in $A$, because this would imply that $|M|\geq 3\cdot 2=6$, so $M$ contains 6 odd vertices, which together with $A$ are 9 odd vertices, contradicting the fact that $Q_{4}$ has only 8 odd vertices in total. It follows that we can add an edge from $a$ to one of the vertices in $A$ without creating a near half-layer, and then the remaining two vertices in $A$ can be matched to each other. ∎

### v2

@@ Abstract.
An extended abstract of this paper appeared in the Proceedings of IWOCA 2024 [FM24].
This work was supported by Czech Science Foundation grant GA 22-15272S. Both authors participated in the workshop ‘Combinatorics, Algorithms and Geometry’ in March 2024, which was funded by German Science Foundation grant 522790373.
@@ 1. Introduction
Figure 1. Cycles and matchings in the hypercube $Q_{4}$: (a) a Hamilton cycle, where the vertex labels omit curly brackets and commas for conciseness; (b) a perfect matching $M$ and an extension of $M$ to a Hamilton cycle; (c) a maximal, but not perfect matching $M^{\prime}$; (d) an extension of $M^{\prime}$ to a cycle factor with two cycles; (e) an extension of $M^{\prime}$ to a (non-Hamilton) cycle; (f) an extension of $M^{\prime}$ to a Hamilton cycle, obtained by joining the two cycles in (d).
@@
In fact, the strengthening to consider $K(Q_{d})$ instead of $Q_{d}$ is the key idea to Fink’s proof, as it makes the induction hypothesis stronger and thus more flexible. This construction actually shows that every perfect matching of $K(Q_{d})$ can be extended to at least $2^{2^{d-4}}$ distinct Hamilton cycles. This proof technique has been exploited in several subsequent papers (see e.g. [Gre09, GNŠ18, AAA+15]), and we will also use it heavily in our arguments.
@@ 1.1. Our results
Furthermore, we can give some guarantee about the length of the cycle obtained from our construction, namely that it visits a constant fraction of the vertex set $V(Q_{d})$. This follows by using that any matching can be extended to a matching that contains a constant fraction of all vertices.
@@
To prove Theorem 5 inductively, we use the following auxiliary theorem, which yields a cycle that extends a given matching but also avoids one ‘forbidden’ vertex $z$, i.e., avoiding $z$ is an additional constraint imposed on the cycle. Using that $Q_{d}$ is vertex-transitive, we may assume w.l.o.g. that the forbidden vertex is $z=\emptyset$. The following theorem requires some additional conditions, and to state them we need to introduce some notation: We split $Q_{d}$ in a direction $i\in[d]$ into subgraphs $Q^{i}_{0}$ and $Q^{i}_{1}$, where $Q^{i}_{1}$ is the subgraph induced by vertices of $Q_{d}$ containing $i$ and $Q^{i}_{0}$ is the subgraph induced by vertices of $Q_{d}$ not containing $i$. Note that $z\in V(Q^{i}_{0})$ for every $i\in[d]$. The set of edges of $Q_{d}$ that have one end vertex in $Q^{i}_{0}$ and the other in $Q^{i}_{1}$ is called a layer in direction $i$. This set splits into two equal-sized sets of edges, depending on the parity of their end vertices in $Q^{i}_{0}$, and we refer to each of these two sets of edges as a half-layer. In this way, all $d2^{d-1}$ edges of $Q_{d}$ are partitioned into $d$ many layers each of cardinality $2^{d-1}$, or $2d$ many half-layers, each of cardinality $2^{d-2}$.
@@
Let $d\geq 5$. A matching $M$ of $K(Q_{d})$ that avoids $z=\emptyset$ can be extended to a cycle that avoids $z$ if and only if it satisfies the following condition:
@@
We remark that the statement of Theorem 8 is true for $d=2$ and $d=3$, but false for $d=4$, and all counterexamples can be seen in [GNŠ18, Figs. 7–9].
@@ 1.2. Applications to Hamilton-laceability
Let $d\geq 5$, and let $x,y$ be two vertices of opposite parity in $Q_{d}$. A perfect matching $M$ of $B(Q_{d})$ with $xy\notin M$ can be extended to a Hamilton path with end vertices $x$ and $y$ if and only if $(M\setminus\{xx^{M},yy^{M}\})\cup\{x^{M}y^{M}\}$ contains no half-layers.
@@
In this theorem, $Q_{d}\setminus\{x,y\}$ represents the graph obtained from $Q_{d}$ by removing the two vertices $x$ and $y$.
@@
Let $d\geq 5$, and let $x,y$ be two vertices of opposite parity in $Q_{d}$. A perfect matching $M$ of $K(Q_{d})$ with $xy\notin M$ can be extended to a Hamilton path with end vertices $x$ and $y$ if and only if $(M\setminus\{xx^{M},yy^{M}\})\cup\{x^{M}y^{M}\}$ contains no half-layers.
@@ 1.3. Avoiding and including other structures
Instead of forbidding edges, one may also consider the problem of forbidding certain vertices. To this end, Fink and Gregor [FG12] proved that for any set $F$ of at most $\binom{d}{2}-2$ vertices of $Q_{d}$, $d\geq 3$, there is a cycle of length at least $2^{d}-2|F|$ that avoids all vertices in $F$; see Figure 2 (d). Their result answers a conjecture by Castañeda and Gotchev [CG10], and the bound $\binom{d}{2}-2$ is best possible. Note that if we forbid some number of vertices of the same parity, then any extending cycle will also skip the same number of vertices of the opposite parity. This explains that we need to allow skipping $2|F|$ many vertices instead of only $|F|$.
@@ 1.4. Other restricted Gray codes
Hamilton cycles and paths in the hypercube are often referred to as Gray codes [Müt23, Sav97]. In this setting, we view the vertices of the hypercube as bitstrings of length $n$, given by the indicator vectors of the corresponding subsets. The well-known binary reflected Gray code, named after the Bell Labs researcher Frank Gray, is a particularly elegant construction of such a cycle that lends itself well to fast algorithmic computation, optimally so that each bit flipped is computed in constant time. In addition to the work discussed in the preceding section about Gray codes that include or avoid certain substructures, there has been a large amount of work on Gray codes that satisfy various other kinds of constraints, motivated by different applications:
@@
  • single-track Gray codes [Etz07, GMM24], where each bit follows the same cyclically shifted flipping pattern along the code.
@@ 1.5. Outline of this paper
In Section 2 we present some notations and terminology used throughout this paper, as well as some auxiliary lemmas needed later. In Sections 3 and 4 we present the inductive arguments for the proofs of Theorems 5 and 8, respectively. Section 5 contains the proofs of Theorems 6, 7, and 12. We conclude with some open problems in Section 6.
@@ 2.1. Notation and definitions
For two sets $A$ and $B$, the symmetric difference of $A$ and $B$ is denoted by $A\mathbin{\triangle}B$. Note that $Q_{d}$ has an edge between any two vertices $u$ and $v$ with $|u\mathbin{\triangle}v|=1$. For a vertex $u$ of $Q_{d}$ and $i\in[d]$, we define $u^{i}:=u\mathbin{\triangle}\{i\}$, and we refer to $i$ as the direction of the edge $uu^{i}$. We say that $u$ is even or odd, if $|u|$ is even or odd, respectively.
@@
For a graph $G$ and a subset of edges $M$ of $G$, we say that a vertex $u$ of $G$ is covered by $M$, if $u$ is an end vertex of one of the edges of $M$. We write $V(M)$ for the set of all vertices covered by $M$. If $u$ is not covered by $M$ we say that it is avoided by $M$. Recall that for a vertex $u$ covered by an edge of a matching $M$, we write $u^{M}$ for the other end vertex of that edge of $M$. Furthermore, for $i\in[d]$ we use the abbreviation $u^{iM}:=(u^{i})^{M}$.
@@ 2.2. A cycle that avoids two vertices
The first lemma asserts that a cycle that extends an almost perfect matching of $K(Q_{d})$ and avoids a particular vertex $x$ is forced to also avoid another vertex $y$, provided that $x$ and $y$ have the correct parity. This is useful when applying induction in the proof of Theorem 8, because it prevents that the invariant ‘cycle avoids one prescribed vertex’ blows up to ‘cycle avoids two prescribed vertices’, ‘cycle avoids three prescribed vertices’, etc., which would be undesirable.
@@
If $C$ avoids $y$, then $C\setminus M$ is set of pairwise disjoint edges whose end vertices have opposite parity (i.e., a matching), covering all vertices of $K(Q_{d}\setminus\{x,y\})$. Since $Q_{d}$ has the same number of even and odd vertices, it follows that the avoided vertices $x$ and $y$ also have opposite parity. On the other hand, if $C$ contains $y$, then $C\setminus M$ is a set of pairwise disjoint edges whose end vertices have opposite parity, plus a path of length 2 whose two end vertices have opposite parity to the middle vertex $y$. It follows that $x$ must have the same parity as $y$. ∎
@@ 2.3. Properties of half-layers and quad-layers
To prove (i), let $i$ and $j$ be the directions of $L$ and $L^{\prime}$, respectively, and consider the partition of $Q_{d}$ into disjoint 4-cycles of the form $C=(u,u^{i},(u^{i})^{j},u^{j})$ for some $u\in V(Q_{d})$. From the two pairs of opposite edges in directions $i$ and $j$ in $C$, exactly one edge belongs to $L$ and one edge to $L^{\prime}$, respectively.
@@
The next lemma asserts that a matching cannot contain half-layers and quad-layers in several different directions.
@@
To prove (iii) we argue as follows. By Lemma 14 (i), two half-layers in different directions share $2^{d-2}$ vertices. Having two 2-near half-layers in different directions, each one uses $22^{d-2}-$ 2 of the shared vertices to be matched in its direction (and not the other direction). However, $2(2^{d-2}-2)>2^{d-2}$ for $d\geq 5$, so two 2-near half-layers in different directions cannot exist for $d\geq 5$. ∎
@@
Let $d=5$, and let $M$ be a matching of $K(Q_{d})$ that avoids $z=\emptyset$ and contains a $z$-dangerous (near) quad-layer. Then there is a direction $i\in[d]$ such that $|M^{i}_{-}|\geq 3$ and $M^{i}_{0}$ contains no (near) half-layers of $Q^{i}_{0}$.
@@
Let $i_{1}$ be the direction of a $z$-dangerous (near) quad-layer, implying that $|M^{i_{1}}_{-}|\geq 2^{d-3}-1=3$. If $M^{i_{1}}_{0}$ contains no (near) half-layers, then we are done. Otherwise, let $i_{2}$ be the direction of a (near) half-layer of $Q^{i_{1}}_{0}$, implying that $|M^{i_{2}}_{-}|\geq 3$. If $M^{i_{2}}_{0}$ contains no (near) half-layers, then we are done. Otherwise, let $i_{3}$ be the direction of a (near) half-layer of $Q^{i_{2}}_{0}$, implying that $|M^{i_{3}}_{-}|\geq 3$ etc. If this process terminates, then we are done. Otherwise, it loops in a cycle of length $\ell\geq 2$. As each of these $\ell$ many $z$-dangerous (near) quad-layers of $Q_{d}$ in direction $i_{j}$, $j\in\{1,\ldots,\ell\}$, covers at least $d-3=2$ neighbors of $z$ with an edge in direction $i_{j}$, we conclude that $\ell=2$. For $\ell=2$, we have a (near) half-layer of $Q^{i_{1}}_{0}$ in direction $i_{2}$ and a (near) half-layer of $Q^{i_{2}}_{0}$ in direction $i_{1}$, and each of them covers at least $d-3=2$ neighbors of $z$ in the subcube $(Q^{i_{1}}_{0})^{i_{2}}_{0}$ with an edge in direction $i_{2}$ and $i_{1}$, respectively. However, the vertex $z$ has only three neighbors in the 3-dimensional subcube $(Q^{i_{1}}_{0})^{i_{2}}_{0}$. ∎
The next lemma allows us to complete partial matchings of $K(Q_{d})$ to perfect matchings without creating half-layers.
@@
We first prove the statement for half-layers. We repeatedly match pairs of vertices in $A$, adding the matched pairs to $P$, thus decreasing the size of $A$ in each round, as follows: If $|A|\geq 6$, we match any two vertices of $A$ having the same parity, which does not create half-layers, and repeat. If $|A|=4$ we distinguish three cases. If all vertices in $A$ have the same parity, we match them arbitrarily, which does not create half-layers. If $A$ has two vertices of each parity, we match them in those pairs, and we are done as well. Otherwise we have one even and three odd vertices in $A$ or vice versa, and by symmetry it suffices to consider the first case. Let $a$ be the unique even vertex in $A$. If one of the three odd vertices in $A$ is not a neighbor of $a$ in $Q_{d}$, then we match $a$ to this vertex and the remaining two vertices to each other, which does not create half-layers. On the other hand, if all three odd vertices in $A$ are neighbors of $a$, then for at most two of them adding an edge to $a$ may create a half-layer. Indeed, if for all three of them adding an edge to $a$ would create a half-layer, then this would imply $|M|\geq 3(2^{d-2}-1)=:\ell$, so $M$ contains $\ell$ odd vertices, which together with $A$ are $\ell+3\geq 3\cdot 2^{d-2}>2^{d-1}$ odd vertices, but $Q_{d}$ has only $2^{d-1}$ odd vertices in total. It follows that we can add an edge from $a$ to one of the vertices in $A$ without creating a half-layer, and then the remaining two vertices in $A$ can be matched to each other.
The same proof works for near half-layers instead of half-layers, the only difference being the inequalities $|M|\geq 3(2^{d-2}-2)=:\ell$ and $\ell+3\geq 3(2^{d-2}-1)>2^{d-1}$ in the last step that are valid for $d\geq 4$. ∎